# Solving the SU($N$) Orbital Hatsugai-Komoto Model

### v2

@@ Abstract
We show that the physics of the SU($N$) Hubbard model can be realistically simulated with the recent advance of the orbital Hatsugai-Kohmoto model. In this approach, the momentum mixing absent from the band Hatsugai-Kohmoto model is included by grouping $n-$Hubbard atoms into a cluster. We take advantage of the rapid convergence of this scheme ($1/n^{2}$) and show that for $n\approx O(10)$, quantitative agreement with determinantal quantum Monte Carlo arises for the double occupancy and the compressibility across the Mott transition, as well as the spin structure factor in the Mott insulating state. Additional features of this work are that we can obtain low-temperature physics, which could be qualitatively different from its high-temperature counterpart, and dynamical quantities without resorting to analytical continuation, thereby establishing a clear advantage of the orbital Hatsugai-Kohmoto model over quantum Monte Carlo methodology in the study of strong correlations.
preprint: APS/123-QED
@@ I Introduction
While SU($2$) is the standard group relevant for describing electronic matter, ultra-cold alkaline earth atoms Gorshkov et al. (2010); Cazalilla and Rey (2014) can realize SU($N$) symmetry by having nuclear spins as large as 10 and in some cases 36 in Na40Schindewolf et al. (2022). Consequently, the more general group SU($N$Honerkamp and Hofstetter (2004); Schindewolf et al. (2022); Wu (2006); Cazalilla et al. (2009); Gorshkov et al. (2010); Taie et al. (2012); Cazalilla and Rey (2014); Hofrichter et al. (2016); Ozawa et al. (2018); Wang et al. (2019); Taie et al. (2022); Tusi et al. (2022); Ibarra-García-Padilla and Choudhury (2024); Pasqualetti et al. (2024) is highly relevant to the cold-atom community. Here SU($N$) acts globally transforming any two local eigenstate $|\alpha,i\rangle$ and $|\beta,i\rangle$ via the unitary matrix $U_{\rm\alpha,\beta}$, such that $|\alpha,i\rangle=\sum_{\beta}U_{\alpha\beta}|\beta,i\rangle$. If the interactions are the same for all $N$ spin components, realizing flavor-selective Mott physics is possible regardless of the underlying lattice. Such a realization obviates the complications introduced by multiple orbitals in traditional solid-state systems that exhibit orbital-selective Mott physics de’ Medici et al. (2014); Yi et al. (2015). In terms of magnetism, SU($N$) Hubbard/Heisenberg models Hermele et al. (2009); Tóth et al. (2010); Bauer et al. (2012); Hermele and Gurarie (2011); Romen and Läuchli (2020); Nataf and Mila (2014); Sotnikov and Hofstetter (2014); Sotnikov (2015); Hafez-Torbati and Hofstetter (2018, 2019); Feng et al. (2023); Ibarra-García-Padilla et al. (2023) provide a richer playground than does SU($2$). While a spin pattern with a period of $N$ is the most generic ordering in SU($N$), stripe phases also obtain, for example, a dimerized pattern not formed out of singlets of SU($4$) but rather a 6-dimensional irreducible representation in SU($4$Corboz et al. (2011). That is, the stripe pattern is not composed of ladders of dimers forming singlets, but rather a set of gapless excitations emerge from the breaking of the SU($4$) despite the accompanying reduction in the translational symmetry Corboz et al. (2011). SU($N$) physics is not limited to single atomic nuclei and extends to, for example, molecular systems as well. Recent experiments Schindewolf et al. (2022) demonstrate that because $s$-wave collisions of shielded molecules have a limited dependence on the underlying spin states, such systems provide a perfect playground for exploring SU($N$) physics. Although the interaction between such molecules is of the standard di-polar form, the sign and magnitude of the scattering length can be tuned by the details of the shielding field. As a result, shielded molecular aggregates from atomic nuclei can even exhibit bosonic or fermionic statistics.
@@
between two dipolar molecules of magnitude $\mu_{1}$ and $\mu_{2}$ separated by a distance $R$. For $s$-wave scattering, although this interaction averages to zero for total angular momentum $L=0$, matrix elements connnect the $L=0$ and $L=2$ states that have both diagonal and off-diagonal components. The off-diagonal component is attractive and scales Yi and You (2000) as $d^{4}/R^{4}$ where $d$ is the space-fixed dipole moment of each molecule. The repulsive interaction arises specifically from the energy shift, $\Delta E$, induced by the external static-field and microwave shielding fields. The matrix element connecting pair states is proportional to $d^{4}/(\Delta ER^{6})$ and hence is shorter range relative to the attractive interaction. It is the interplay between the short-range repulsion and long-range attraction that generates confined and deconfined phases that are color specific Mukherjee et al. (2025).
While the knee-jerk instinct is that SU($N$) in the large $N$ limit is inherently mean-field, this is not the case for describing the details of the phase diagram for ultra-cold atoms. Indeed, discerning the phase diagram of SU($N$) matter is particularly taxing computationally, as the size of Hilbert space scales as $2^{N\times N_{\rm sites}}$. Consequently, advancing the physics of SU($N$) systems requires new tools. In this study, we adopt the newly developed orbital Hatsugai-Kohmoto (OHK) model Mai et al. (2024). This model builds in momentum mixing into the tractable HK model Hatsugai and Kohmoto (1992); Phillips et al. (2020); Huang et al. (2022) (now referred to as band HK). If $n$ is the number of momenta that are mixed, the convergence to Hubbard scales as $1/n^{2}$. Such rapid convergence makes this a tractable scheme for simulating Hubbard physics. It is this scheme that we employ here for the SU($N$) problem. We compute the double occupancy, compressibility, Mott gap and the dynamical spin structure factor across the Mott transition. We obtain the latter directly without the need for analytical continuation as is required in quantum Monte Carlo methods. Our results show quantitative agreement with these numerical approaches where comparison is possible, and we extend our analysis to regions where these methods become impractical.
@@
Figure 1: Panels (a-d) show density $\langle\hat{n}\rangle$ and compressibility $\chi=\partial\langle\hat{n}\rangle/\partial\mu$ versus chemical potential $\mu$ with SU($N$) 4-OHK in panels (a,c) and SU($N$) Hubbard obtained with DQMC Ibarra-García-Padilla et al. (2021) in panels (b,d). $\mu_{0}$ is calculated such that $\langle\hat{n}\rangle|_{\mu_{0}}=1$. We fix $U=8t$ and $\beta=2t^{-1}$, and find $\mu_{0}/U=0.49,0.40,0.35$ for $N=2,3,4$, respectively. Panels (e) and (f) show double occupancy $\mathcal{D}$ and its derivative $\partial\mathcal{D}/\partial\langle\hat{n}\rangle$ versus $\langle\hat{n}\rangle$, respectively. Here we fix $U=8t,\ \beta=2t^{-1}$. The line and scatter plots represent 4-OHK and DQMC-Hubbard results Ibarra-García-Padilla et al. (2021), respectively.
@@ II Model and methods
The SU($N$) generalization of the band HK model is exactly solvable and was introduced soon after the original band HK model Nogueira and Anda (1996) (see supplement for a discussion of this model). However, the kinetic and interaction terms commute with each other in the SU($N$) band HK model, meaning quantum fluctuations are not captured. Here, we introduce the SU($N$) generalization of the OHK model, where quantum fluctuations and nontrivial dynamics are incorporated through momentum mixing Mai et al. (2024):
@@
$H_{\text{OHK}}^{N}=$ $\sum_{\alpha\alpha^{\prime}\sigma,{\bf k}\in{\rm rBZ}}\xi^{\alpha\alpha^{\prime}}_{\bf k}c^{\dagger}_{\alpha{\bf{\bf k}}\sigma}c_{\alpha^{\prime}{\bf k}\sigma}$
$+\frac{U}{2}\sum_{\alpha,{\bf k}\in{\rm rBZ}}\sum_{\sigma\neq\sigma^{\prime}}n_{\alpha{\bf k}\sigma}n_{\alpha{\bf k}\sigma^{\prime}},$ (2)
where $\alpha,\alpha^{\prime}=1,...,n$ denote orbital indices and $\sigma,\sigma^{\prime}=1,...,N$ denote the flavor degree of freedom. The orbital degrees of freedom are introduced by real space clustering of atoms, i.e. choice of unit cell, and $\xi^{\alpha\beta}_{\bf k}$ is the resulting hopping matrix. We work on a $2D$ square lattice, where in the $n=1$ band HK limit, the hopping matrix reduces to the familiar $\xi_{\bf k}=-2t(\cos{k_{x}}+\cos{k_{y}})-\mu$, where $t=1$ is the nearest neighbour hopping setting the energy scale, $\mu$ is the chemical potential, and momentum ranges over the original Brillouin zone (BZ) $k_{x},k_{y}\in[0,2\pi)$. For $n>1$, $\xi_{\bf k}$ turns into $\xi^{\alpha\alpha^{\prime}}_{\bf k}$ and the momentum summation is restricted to a reduced BZ (rBZ) determined by the particular choice of unit cell, thereby ensuring that the kinetic term remains unchanged. Contrastly, the interaction term formally retains the same structure but effectively depends on the orbital choice. Notably, translational invariance in the Hubbard model is preserved in the OHK formulation. This follows from the fact that the summation over $\bf k$ in Eq. (II) can be extended to the full Brillouin zone, and introducing an overall factor of $1/n$ leaves the Hamiltonian unchanged. We choose our clustering to preserve the $C_{4}$ symmetry of the lattice, unless specified otherwise. Note that for $N=2$, this model reduces to the SU($2$) OHK model recently introduced in Ref. Mai et al. (2024). Hubbard physics obtains exactly in the $n\rightarrow\infty$ limit. However, since the convergence to Hubbard scales as $1/n^{2}$, $n$ needs not to be excessively large for quantitative agreement with the Hubbard model.
@@
All numerical results presented in this work are obtained using exact diagonalization of the Hamiltonian in Eq. (II). We obtain dynamical quantities by applying the Lanczos method to the ground state obtained from exact diagonalization. The 4-OHK calculations are simulated on the rBZ with $L\times L$ points where $L=20$. For 8- and 9-OHK calculations, we set $L=10$.
@@
Figure 2: Panel (a) shows the density $\langle\hat{n}\rangle$ versus chemical potential $\mu$ at $U=8t$ and $\beta=200t^{-1}$ for varying $N=2,3,4$ in SU($N$) 4-OHK models. Panels (b) and (c) show the charge gap $\Delta$ at $\langle\hat{n}\rangle=1$ filling as a function of interaction strength $U$ at zero temperature for varying $N$ in SU($N$) 4- and 8-OHK models, respectively. For the 8-OHK model, we use a $2\times 4$ unit cell.
@@ III Results of SU($N$) OHK
We begin by investigating the generalized SU($N$) OHK models with a focus on their general $N$-dependent behaviors. For a given $N$, Mott insulating states are possible at fillings $\langle\hat{n}\rangle=a$ for $a=1,...,N$. Of particular interest are the cases of one particle per site, $\langle\hat{n}\rangle=1$, and half-filling, $\langle\hat{n}\rangle=N/2$ (when $N$ is even). These two phases coincide for $N=2$ and hence they are natural SU($N$) generalizations of the SU(2) half-fillied Mott insulator.
III.1 Mott insulator at $\langle\hat{n}\rangle=1$
High temperature. We first solve 4 ($2\times 2$)-OHK at high temperature $\beta=2t^{-1}$ enabling a direct comparison with previous determinantal quantum Monte Carlo (DQMC) simulations on a $6\times 6$ SU($N$) Hubbard cluster Ibarra-García-Padilla et al. (2021). In Fig. 1, we show the filling $\langle\hat{n}\rangle$ and compressibility $\chi\equiv\partial\langle\hat{n}\rangle/\partial\mu$ as a function of $\mu$ for the 4-OHK model in panels (a) and (c), and the corresponding DQMC results are shown in panels (b) and (d). We already find qualitative agreement between 4-OHK and the state-of-the-art DQMC simulations on the Hubbard model. This is expected given the rapid convergence of OHK to the Hubbard model. In particular, we observe the softening of the Mott gap as $N$ increases in both the density and compressibility. However, there are quantitative differences, for example, 4-OHK overestimates the magnitude of the Mott gap.
In Fig. 1(e,f), we present the benchmark for double occupancy $\mathcal{D}$, and its derivative $\partial\mathcal{D}/\partial\langle\hat{n}\rangle$ as a function of density $\langle\hat{n}\rangle$. Double occupancy in the OHK model is defined as

### v3

# Solving the SU($N$) Orbital Hatsugai-Kohmoto Model

@@ Abstract
We show that the physics of the SU($N$) Hubbard model can be realistically simulated with the recently developed orbital Hatsugai-Kohmoto model. In this approach, the momentum mixing absent from the band Hatsugai-Kohmoto model is included by grouping $n$-Hubbard atoms into a cluster. We take advantage of the rapid convergence of this scheme ($1/n^{2}$) and show that for $n\approx O(10)$, quantitative agreement with the determinantal quantum Monte Carlo method arises for the double occupancy as well as qualitative agreement for the filling and compressibility across the Mott transition. Additional features of this work are that we can obtain low-temperature physics, which could be qualitatively different from its high-temperature counterpart, and dynamical quantities without resorting to analytical continuation, thereby establishing the orbital Hatsugai-Kohmoto model as a useful alternative perspective for studying strong correlations in SU($N$) systems.
††preprint: APS/123-QED
@@ I Introduction
While SU($2$) is the standard group relevant for describing electronic matter, ultra-cold alkaline earth atoms [1, 2] can realize SU($N$) symmetry by having nuclear spins as large as 10 and in some cases 36 in Na40K [3]. Consequently, the more general group SU($N$) [4, 3, 5, 6, 1, 7, 2, 8, 9, 10, 11, 12, 13, 14] is highly relevant to the cold-atom community. Here SU($N$) acts globally transforming any two local eigenstates $|\alpha,i\rangle$ and $|\beta,i\rangle$ via the unitary matrix $U_{\rm\alpha,\beta}$, such that $|\alpha,i\rangle=\sum_{\beta}U_{\alpha\beta}|\beta,i\rangle$. If the interactions are the same for all $N$ spin components, realizing flavor-selective Mott physics is possible regardless of the underlying lattice. Such a realization obviates the complications introduced by multiple orbitals in traditional solid-state systems that exhibit orbital-selective Mott physics [15, 16]. In terms of magnetism, SU($N$) Hubbard and Heisenberg models [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] provide a richer playground than does SU($2$). While a spin pattern with a period of $N$ is the most generic ordering in SU($N$), stripe phases also obtain, for example, a dimerized pattern not formed out of singlets of SU($4$) but rather a six-dimensional irreducible representation in SU($4$) [29]. That is, the stripe pattern is not composed of ladders of dimers forming singlets, but rather a set of gapless excitations emerge from the breaking of the SU($4$) despite the accompanying reduction in the translational symmetry [29]. SU($N$) physics is not limited to single atomic nuclei and extends, for example, to molecular systems as well. Recent experiments [3] demonstrate that because $s$-wave collisions of shielded molecules have a limited dependence on the underlying spin states, such systems provide a perfect playground for exploring SU($N$) physics. Although the interaction between such molecules is of the standard dipolar form, the sign and magnitude of the scattering length can be tuned by the details of the shielding field. As a result, shielded molecular aggregates from atomic nuclei can even exhibit bosonic or fermionic statistics.
@@
between two dipolar molecules of magnitude $\mu_{1}$ and $\mu_{2}$ separated by a distance $R$. For $s$-wave scattering, although this interaction averages to zero for total angular momentum $L=0$, matrix elements connect the $L=0$ and $L=2$ states that have both diagonal and off-diagonal components. The off-diagonal component is attractive and scales [30] as $d^{4}/R^{4}$ where $d$ is the space-fixed dipole moment of each molecule. The repulsive interaction arises specifically from the energy shift, $\Delta E$, induced by the external static field and microwave shielding fields. The matrix element connecting pair states is proportional to $d^{4}/(\Delta ER^{6})$ and hence is shorter range relative to the attractive interaction. It is the interplay between the short-range repulsion and long-range attraction that generates confined and deconfined phases that are color specific [31].
While the knee-jerk instinct is that SU($N$) in the large $N$ limit is inherently mean-field, this is not the case for describing the details of the phase diagram for ultra-cold atoms. Indeed, discerning the phase diagram of SU($N$) matter is particularly taxing computationally, as the size of the Hilbert space scales as $2^{N\times N_{\rm sites}}$. Consequently, advancing the physics of SU($N$) systems requires new tools. In this study, we adopt the recent orbital Hatsugai-Kohmoto (OHK) model [32, 33]. This model builds in momentum mixing into the tractable HK model [34, 35, 36] (now referred to as band HK or $1$-OHK). If $n$ is the number of momenta that are mixed, the convergence to Hubbard scales as $1/n^{2}$. Such rapid convergence makes this a tractable scheme for simulating Hubbard physics. It is this scheme that we employ here for the SU($N$) problem. We compute the double occupancy, compressibility, Mott gap, and dynamical spin structure factor across the Mott transition. We obtain the latter directly from exact diagonalization without the need for analytical continuation as is required in quantum Monte Carlo methods. Our results show quantitative agreement with these numerical approaches where comparison is possible, and we extend our analysis to regions where these methods become impractical.
@@
Figure 1: Panel (a) is a schematic showing the real-space clustering for 4-OHK and some representative elements of the hopping matrix $\xi^{\alpha\alpha^{\prime}}_{\mathbf{k}}$. Panel (b) shows the original BZ, i.e. the $n=1$ limit, and the rBZ4 for 4-OHK (for a $2\times 2$ unit cell).
[…]
Figure 2: Panels (a-d) show density ⟨$\langle\hat{n}\rangle$ and compressibility $\chi=\partial\langle\hat{n}\rangle/\partial\mu$ versus chemical potential $\mu$ with SU($N$) 4-OHK in panels (a,c) and SU($N$) Hubbard obtained with DQMC [37] in panels (b,d). $\mu_{0}$ is calculated such that ⟨$\langle\hat{n}\rangle|_{\mu_{0}}=1$. We fix $U=8t$ and $\beta=2t^{-1}$, and find $\mu_{0}/U=0.49,0.40,0.35$ for $N=2,3,4$, respectively. Panels (e) and (f) show double occupancy $\mathcal{D}$ and its derivative ∂$\partial\mathcal{D}/\partial\langle\hat{n}\rangle$ versus ⟨$\langle\hat{n}\rangle$, respectively. Here we fix $U=8t,\ \beta=2t^{-1}$. The line and scatter plots represent 4-OHK and DQMC-Hubbard results [37], respectively.
@@ II Model and methods
This work aims to establish OHK models as an efficient tool for simulating SU($N$) Hubbard physics. We begin with a review of the connection between the Hubbard model and OHK models following the discussion in Ref. [33]. Consider first the Hubbard interaction
$H_{\text{HB}}^{\text{int}}$ $=U\sum_{i}n_{\mathbf{r}_{i}\uparrow}n_{\mathbf{r}_{i}\downarrow}$
$=\frac{U}{N_{s}}\sum_{\mathbf{k,p,q}\in\text{BZ}}c^{\dagger}_{{\bf k}\uparrow}c_{{\bf k}-\bf q\uparrow}c^{\dagger}_{{\bf k}+{\bf p}\downarrow}c_{{\bf k}+{\bf p}+{\bf q}\downarrow},$ (2)
where $N_{s}$ is the number of sites, and the band HK interaction
$H^{\text{int}}_{\text{HK}}=U\sum_{\mathbf{k\in\text{BZ}}}c^{\dagger}_{\mathbf{k}\uparrow}c_{\mathbf{k}\uparrow}c^{\dagger}_{\mathbf{k}\downarrow}c_{\mathbf{k}\downarrow}.$ (3)
Note that we restrict our attention to SU(2) for simplicity. The momentum space representation of the Hubbard interaction highlights the deviation between the two models, namely, band HK contains no momentum scattering. We can move towards Hubbard physics by systematically introducing scattering momenta ${\bf p,q}\in B_{n}$ to the interaction term, leading to the OHK model. We refer to the OHK model with $n$ scattering momenta (or equivalently, $n$ orbitals, as illustrated below) as the $n$-OHK model. The case of $n=4$ is depicted in Fig. 1. Here we set $B_{4}=\{(0,0),(0,\pi),(\pi,0),(\pi,\pi)\}$ and consider an OHK interaction
$H^{\text{int},n=4}_{\text{OHK}}=\frac{U}{4}\sum_{\mathbf{k}\in\text{BZ}}\sum_{\mathbf{p,q}\in B_{4}}c^{\dagger}_{{\bf k}\uparrow}c_{{\bf k}-\bf q\uparrow}c^{\dagger}_{{\bf k}+{\bf p}\downarrow}c_{{\bf k}+{\bf p}+{\bf q}\downarrow}.$ (4)
The corresponding real-space picture is shown in Fig. 1(a), where each unit cell includes four sites/orbitals. We can consider a transformation to the orbital basis $c_{\alpha\mathbf{k}\sigma}=\frac{1}{2}\sum_{\mathbf{q}\in B_{4}}c_{\mathbf{k+q}\sigma}e^{-i\mathbf{q}\cdot\mathbf{r_{\alpha}}}$, where $\alpha=1,...,4$ labels the orbitals and $\bf r_{\alpha}$ denotes the position of each orbital within the unit cell. Performing this transformation on Eq. 4 leads to
$H^{\text{int},n=4}_{\text{OHK}}=U\sum_{\alpha=1}^{4}\sum_{\mathbf{k}\in\text{rBZ}_{4}}c^{\dagger}_{\alpha\bf k\uparrow}c_{\alpha\bf k\uparrow}c^{\dagger}_{\alpha\bf k\downarrow}c_{\alpha\bf k\downarrow},$ (5)
where the momentum summation is restricted to a reduced Brillouin Zone (rBZn), reflecting the enlarged unit cell in real space (see Fig. 1(b)). We see that the interaction term reduces to a form similar to Eq. 3, however, with additional orbital degrees of freedom. This completes the construction of the interaction term of the 4-OHK model. This construction can be generalized to an arbitrary number of scattering momenta, and Hubbard physics in the thermodynamic limit is guaranteed to be obtained as $n\rightarrow\infty$ [33]. The $n^{2}$ terms in Eq. 4 introduced by $n$ scattering momenta, along with the numerical evidence from [33], indicate that the convergence to Hubbard scales at least as fast as $1/n^{2}$. Therefore, $n$ does not need to be excessively large for quantitative agreement with the Hubbard model. Note that while the addition of momentum scattering modifies the interaction, we ensure that the kinetic term in the Hamiltonian is kept unchanged throughout this process
@@
∑$\sum_{\bf k\in\text{BZ},\sigma}\xi_{\bf k}c^{\dagger}_{\bf k\sigma}c_{\bf k\sigma}=\sum_{\begin{subarray}{c}\mathbf{k}\in{\rm rBZ}_{n},\\ \alpha\alpha^{\prime},\sigma\end{subarray}}\xi^{\alpha\alpha^{\prime}}_{\bf k}c^{\dagger}_{\alpha\bf k\sigma}c_{\alpha^{\prime}\bf k\sigma}.$ (6)
Here $\xi^{\alpha\alpha^{\prime}}_{\bf k}=\sum_{\delta\mathbf{r}}t^{\alpha\alpha^{\prime}}_{\delta\mathbf{r}}e^{i\mathbf{k}\cdot(\mathbf{\delta r+r_{\alpha}-r_{\alpha^{\prime}}})}-\mu\delta_{\alpha,\alpha^{\prime}}$, where $t^{\alpha\alpha^{\prime}}_{\delta\mathbf{r}}$ is the hopping between $\alpha$ and $\alpha^{\prime}$ orbitals in unit cells separated by $\delta\mathbf{r}$, and $\mu$ is the chemical potential. In this work, we consider a two-dimensional square lattice with nearest-neighbour hopping, where $\xi_{\bf k}=-2t(\cos{k_{x}}+\cos{k_{y}})-\mu$, with $t=1$ setting the energy scale. Representative off-diagonal elements of $\xi^{\alpha\alpha^{\prime}}_{\bf k}$ for the 4-OHK model are shown schematically in Fig. 1(a). Note that, as a result of the orbital mixing, the kinetic and potential terms no longer commute. It is this new ingredient that introduces spectral weight transfer on the Mott scale.
The SU($N$) generalization[38] of the band HK model is exactly solvable and was introduced soon after the original band HK model(see the appendix for a discussion of this model). The introduction of momentum mixing to the SU($N$) HK model follows straightforwardly from the SU(2) case, allowing us to introduce the SU($N$) OHK model
$H_{\text{OHK}}^{N}=$$\sum_{\begin{subarray}{c}{\bf k}\in{\rm rBZ}_{n},\\ \alpha\alpha^{\prime},\sigma\end{subarray}}\xi^{\alpha\alpha^{\prime}}_{\bf k}c^{\dagger}_{\alpha{\bf{\bf k}}\sigma}c_{\alpha^{\prime}{\bf k}\sigma}$
$+\frac{U}{2}\sum_{\begin{subarray}{c}{\bf k}\in{\rm rBZ}_{n},\\ \alpha,\sigma\neq\sigma^{\prime}\end{subarray}}n_{\alpha{\bf k}\sigma}n_{\alpha{\bf k}\sigma^{\prime}},$ (7)
where $\alpha,\alpha^{\prime}=1,...,n$ denote orbital indices and $\sigma,\sigma^{\prime}=1,...,N$ denote the flavor degree of freedom. In the case of SU($N$), the inclusion of $n$ orbitals introduces $N(N-1)n^{2}/2$ terms when expressing the interaction in the form of Eq. 4. This suggests faster convergence to Hubbard physics with increasing $n$ when compared to the SU(2) case.
@@
All numerical results presented in this work are obtained using exact diagonalization of the Hamiltonian in Eq. II. We obtain dynamical quantities by applying the Lanczos method to the ground state obtained from exact diagonalization. The 4-OHK calculations are simulated on the rBZ4 with $L\times L$ points where $L=20$. For 8- and 9-OHK calculations, we set $L=10$.
@@
Figure 3: Panel (a) shows the density ⟨$\langle\hat{n}\rangle$ versus chemical potential $\mu$ at $U=8t$ and $\beta=200t^{-1}$ for varying $N=2,3,4$ in SU($N$) 4-OHK models. Panels (b) and (c) show the charge gap $\Delta$ at ⟨$\langle\hat{n}\rangle=1$ filling as a function of interaction strength $U$ at zero temperature for varying $N$ in SU($N$) 4- and 8-OHK models, respectively. For the 8-OHK model, we use a $2\times 4$ unit cell.
@@ III Results of SU($N$) OHK
We begin by investigating the generalized SU($N$) OHK models with a focus on their general $N$-dependent behaviors. For a given $N$, Mott insulating states are possible at fillings ⟨$\langle\hat{n}\rangle=a$ for $a=1,...,N$. Of particular interest are the cases of one particle per site, ⟨$\langle\hat{n}\rangle=1$, and half-filling, ⟨$\langle\hat{n}\rangle=N/2$ (when $N$ is even). These two phases coincide for $N=2$ and hence they are natural SU($N$) generalizations of the SU(2) half-filled Mott insulator.
III.1 Mott insulator at ⟨$\langle\hat{n}\rangle=1$
High temperature. We first solve 4-OHK (using a $2\times 2$ unit cell), at high temperature $\beta=2t^{-1}$ enabling a direct comparison with previous determinantal quantum Monte Carlo (DQMC) simulations on a $6\times 6$ SU($N$) Hubbard cluster [37]. In Fig. 2, we show the filling